OpenStudy (anonymous):

help please? Becky purchased a home entertainment center set for $2,254 using an 18-month deferred payment plan with an interest rate of 22.48%. She did not make any payments during the deferment period. What will Becky’s monthly payment be if she must pay off the dining room set within three years after the deferment period?

OpenStudy (anonymous):

@surry99 could you help me with this one too? i want to make sure i'm doing them right.

OpenStudy (surry99):

Sure you get started and I will be right back...helping someone else at the moment

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

2,254(.2248)=506.70 2,254+506.70=2,760.70 2,760.70(.2248)=620;61 506.70+620.61+2,254=3,381.31 is my final answer i believe

OpenStudy (anonymous):

not sure if i did that right or not

OpenStudy (surry99):

you have done the first two years you need to do the third year

OpenStudy (anonymous):

so what i have so far is correct? what should i do now?

OpenStudy (anonymous):

in order to get the third year

OpenStudy (surry99):

Just continue your calculations from 3381.31

OpenStudy (anonymous):

should i multiply that from the interest? then add that to 3381.31 to get the final answer?

OpenStudy (anonymous):

2,254(.2248)=506.70 2,254+506.70=2,760.70 2,760.70(.2248)=620;61 506.70+620.61+2,254=3,381.31 3,381.31(.2248)=760.12 3,381.31+760.12=4,141.43 as my final answer.

OpenStudy (surry99):

Looks good ...so she owes a total of 4141.43 and must pay that off in how many months?

OpenStudy (surry99):

oops hang on..

OpenStudy (surry99):

For first year 2254(.2248) = 506.70 so at the end of first year...2254 + 506.70 = 2760.70 Second year 2760.70(.2248) =620.61 end of second year 2760.70+ 620.67 = 3381.31 Third year 3381.31(.2248) = 760.12 End of third year = 3381.31+760.12 = 4141.43 ok, just writing it a little different. Now how many months does she have to pay this total amount off...Read the question carefully

OpenStudy (anonymous):

she has 3 years

OpenStudy (anonymous):

or wait is it 18 months?

OpenStudy (surry99):

no she has eighteen months because she makes no payments during the deferment period

OpenStudy (anonymous):

okay so how do i figure out her monthly payments for those 18 months?

OpenStudy (surry99):

Take the total she owes and divide by 18 months...make sense?

OpenStudy (anonymous):

4141.43/18=230.08

OpenStudy (surry99):

looks good...the question is a bit ambiguous but we assumed the 3 years started from the moment she bought the entertainment center which is what happens in real life.

OpenStudy (anonymous):

thanks so much! you're a great help.

OpenStudy (surry99):

np